### v2

@@ 7 APPENDIX B: Distributed Attack Detection Scheduling Algorithm
The ADS algorithm in literature Suo et al. (2024c) can be extended to the D-ADS algorithm, as shown in Algorithm 3.3. The main difference between the D-ADS and ADS algorithms lies in the sensor selection method, specifically whether it uses centralized selection or distributed two-stage sensor selection. For the $l$-th selection at time step $k$, a suspicious sensor is first chosen from each subset (Stage 1 of sensor selection) and added to the set $\mathcal{S}^{(l)}_{k}$. Then, a sensor is randomly selected from $\mathcal{S}^{(l)}_{k}$ as the result of the $l$-th selection (Stage 2 of sensor selection). The algorithm terminates when $|\mathcal{A}_{i,k}^{(l)}|=q_{i}$.
@@
It should be noted that in steps 4-7, the gains of all remaining sensors $j\in\mathcal{N}_{i}\backslash\mathcal{A}_{i,k}^{(l-1)}$ are updated. However, with the improved partitioning strategy, only the gains of one subset of sensors need to be updated here. The meanings of the parameters in Algorithm 3.3 are summarized in TABLE $2$Other variable symbols that are not explained here are the same as literature Suo et al. (2024c).
@@
2:Suspicious sensor set $\mathcal{A}_{i,k}$ at time $k$ and $\mathcal{A}_{i,k,g_{i}}=\mathcal{A}_{i,k,g_{i}}^{(l)}$, where $g_{i}=1:m_{i}$, $k=1,2,...,T$.
@@
[…]
@@
[…]
@@
[…]
@@
$\mathcal{A}_{i,k,g_{i}^{s}}^{l_{g_{i}^{s}}}$ The set of suspicious sensors after the $l_{g_{i}^{s}}$-th selection for the $g_{i}^{s}$-th subset at moment $k$
@@
Next, the performance relationship between the D-ADS algorithm in this paper and the ADS algorithm in literature Suo et al. (2024c) is analyzed. Both algorithms aim to select the set of suspicious sensors that have the greatest impact on the objective function (6). Therefore, ideally, their performance is consistent, meaning that $\mathcal{A}_{i,k}=\cup^{m_{i}}_{g_{i}=1}{\mathcal{A}_{i,k,g_{i}}}$.
@@
For $g_{i}\in\{1,...,m_{i}\}$, $l\in\{0,1,...,q_{i}\}$, define $\delta_{l,g_{i}}$ as $\delta_{l,g_{i}}=\sum_{k=1}^{T}{(}\frac{1}{m_{i}}f_{k}(\mathcal{A}_{i,k}^{*})-f_{k}(\mathcal{A}_{i,k,g_{i}}^{(l_{g_{i}})}))$ , where $\mathcal{A}_{i,k,g_{i}}^{(l_{g_{i}})}$ represents the set of suspicious sensors after the $l_{g_{i}}$-th selection of the $g_{i}$-th subset at moment $k$. Then, the relationship in literature Suo et al. (2024c) can be transformed as $\sum^{m_{i}}_{g_{i}=1}\delta_{l+1,g_{i}}-(1-\frac{1}{q_{i}})^{l+1}\sum^{m_{i}}_{g_{i}=1}\delta_{0,g_{i}}\leq\frac{m_{i}}{q_{i}}\sum_{j=1}^{l+1}(1-\frac{1}{q_{i}})^{l+1-j}B_{i}^{(j)}$. The parameter $B_{i}^{(l)}$ is described in detail in the proof. At this time, the Lemma 3.2 in literatureSuo et al. (2024c) is extended to the distributed case.
@@
$\frac{1}{m_{i}}\sum^{m_{i}}_{g_{i}=1}\delta_{l,g_{i}}$ (29)
$\leq$ $\frac{1}{m_{i}}\sum^{T}_{k=1}(\sum^{m_{i}}_{g_{i}=1}(f_{k}(\mathcal{A}_{i,k,g_{i}}^{*})-f_{k}(\mathcal{A}_{i,k,g_{i}}^{(l_{g_{i}})})))$
@@
$\frac{1-q_{i}}{m_{i}}\sum^{m_{i}}_{g_{i}=1}\delta_{l,g_{i}}\leq-\frac{q_{i}}{m_{i}}\left(\sum^{m_{i}}_{g_{i}=1}\delta_{l+1,g_{i}}\right)+B^{(l+1)}_{i},$
@@
$B^{(l)}_{i}=\sum_{j=1}^{q_{i}}\sum_{k=1}^{T}\left(\frac{1}{m_{i}}\sum^{m_{i}}_{g_{i}=1}G_{k,g_{i}}^{(l_{g_{i}})}p_{g_{i},k}^{(l_{g_{i}})}-G_{kj_{kl}^{*}}^{(l)}\right),$
@@
$\sum^{m_{i}}_{g_{i}=1}\delta_{l+1,g_{i}}-\left(1-\frac{1}{q_{i}}\right)\sum^{m_{i}}_{g_{i}=1}\delta_{l,g_{i}}\leq\frac{m_{i}}{q_{i}}B_{i}^{(l+1)}$
@@
For the dynamic attack strategy in the Definition 1, the proposed D-ADS algorithm can ensure that the theoretical lower bound of the average optimization rate of suspicious sensor selection for any subset is $1-1/e$, and the error expectation is bounded.
@@
To prove that the theoretical lower bound of the average optimization rate over the entire time period is $1-1/e$, we need to show that the expectation of the error
@@
is bounded, where $f_{k}\left(\mathcal{A}_{i,k,g_{i}}\right)$ represents the submodular function value for the selected suspicious sensor set in subset $g_{i}$ at time $k$.
Essentially, the expectation of the error represents the expected difference between the objective function value of the selected suspicious sensor set and $1-1/e$ times the optimal suspicious sensor set’s objective function value. Therefore, the larger the objective function value of a subset, the greater the upper bound of the error expectation. Consequently, the subset with the largest objective function value has the highest upper bound for error expectation, as shown in equation (7)
@@
Based on the Lemma 3 from literature Matsuoka et al. (2021) and the definition $\delta_{l}=\sum_{k=1}^{T}\left(f_{k}(\mathcal{A}_{i,k}^{*})-f_{k}(\mathcal{A}_{i,k}^{(l)})\right)$ for $l=1:q_{i}$, we obtain that, $\sum_{k=1}^{T}{f_{k}\left(\mathcal{A}_{i,k}^{*}\right)}-\sum_{k=1}^{T}{f_{k}\left(\mathcal{A}_{i,k}\right)}+(1-(1-\frac{1}{q_{i}})^{q_{i}})\sum_{k=1}^{T}{(f_{k}(\mathcal{A}_{i,k}^{(0)})-f_{k}(\mathcal{A}_{i,k}^{*}))}\leq\delta_{q_{i}}-(1-\frac{1}{q_{i}})^{q_{i}}\delta_{0}\leq\sum^{m_{i}}_{g_{i}=1}\delta_{l,g_{i}}-(1-\frac{1}{q_{i}})^{q_{i}}\sum^{m_{i}}_{g_{i}=1}\delta_{0,g_{i}}$, where the inequality $\delta_{q_{i}}\leq\sum^{m_{i}}_{g_{i}=1}\delta_{l,g_{i}}$ follows from the diminishing marginal returns property of the submodular function, and the initial condition satisfies $\delta_{0}=\sum_{g_{i}=1}^{m_{i}}\delta_{0,g_{i}}$. Combining equation (30) with the above relation, equation (7) can be rewritten as
$(1-\frac{1}{e})\sum_{k=1}^{T}{\frac{1}{m_{i}}f_{k}\left(\mathcal{A}_{i,k}^{*}\right)}-\sum_{k=1}^{T}{\max_{g_{i}}f_{k}(\mathcal{A}_{i,k,g_{i}})}$ (33)
$\leq$ $\frac{m_{i}}{q_{i}}\sum_{j=1}^{q_{i}}(1-\frac{1}{q_{i}})^{q_{i}-j}\frac{B_{i}^{(j)}}{m_{i}}.$
Thus, to prove that equation (7) is bounded, it suffices to show that $\mathbb{E}[B_{i}^{(l)}/m_{i}]$ is bounded
$\mathbb{E}\left[\frac{B_{i}^{(l)}}{m_{i}}\right]$
$=$ $\frac{1}{m_{i}}\sum_{j=1}^{q_{i}}\mathbb{E}\left[\sum_{k=1}^{T}\left(\frac{1}{m_{i}}\sum^{m_{i}}_{g_{i}=1}G_{k,g_{i}}^{\left(l_{g_{i}}\right)}p_{g_{i},k}^{\left(l_{g_{i}}\right)}-G_{kj_{kl}^{*}}^{\left(l\right)}\right)\right]$
@@
$\leq$ $\tilde{\mathcal{O}}\left(\frac{q_{i}}{m_{i}}\sqrt{3T+2\varDelta_{T}}\right),$

### v3

@@ Appendix B APPENDIX B: Distributed Attack Detection Scheduling Algorithm
The ADS algorithm in literature Suo et al. (2024c) can be extended to the D-ADS algorithm, as shown in Algorithm 3.3. The main difference between the D-ADS and ADS algorithms lies in the sensor selection method, specifically whether it uses centralized selection or distributed two-stage sensor selection. For the $l$-th selection at time step $k$, a malicious information is first chosen from each subset (Stage 1 of sensor selection) and added to the set $\mathcal{S}^{(l)}_{k}$. Then, a sensor is randomly selected from $\mathcal{S}^{(l)}_{k}$ as the result of the $l$-th selection (Stage 2 of sensor selection). The algorithm terminates when $|\mathcal{A}_{i,\,k}^{(l)}|=q_{i}$.
@@
It should be noted that in steps 4-7, the gains of all remaining sensors $j\in\mathcal{N}_{i}\backslash\mathcal{A}_{i,\,k}^{(l-1)}$ are updated. However, with the improved partitioning strategy, only the gains of one subset of sensors need to be updated here. The meanings of the parameters in Algorithm 3.3 are summarized in TABLE 6222Other variable symbols that are not explained here are the same as literature Suo et al. (2024c).
@@
2:Attacked sensor set $\mathcal{A}_{i,\,k}$ at time $k$ and $\mathcal{A}_{i,\,k,\,g_{i}}=\mathcal{A}_{i,\,k,\,g_{i}}^{(l)}$, where $g_{i}=1:m_{i}$, k=1,2,,Tk=1,2,...,T.
@@
7:   $\%$ With optimized grouping strategies, only a subset of sensors’ gains needs to be updated here.
@@
14:   Select an element $j^{(l_{g_{i}})}_{g_{i},select}$ based on the distribution vector $p_{g_{i},k}^{(l_{g_{i}})}$. $\%$ Stage 1 sensor selection.
@@
17:  Randomly select an element $j_{g_{i}^{s},\,select}^{(l_{g_{i}^{s}})}$ from $\mathcal{S}_{k}^{(l)}$. $\%$ Stage 2 sensor selection.
@@
$\mathcal{A}_{i,\,k,\,g_{i}^{s}}^{l_{g_{i}^{s}}}$ The set of attacked sensors after the $l_{g_{i}^{s}}$-th selection for the $g_{i}^{s}$-th subset at moment $k$
@@
Next, the performance relationship between the D-ADS algorithm in this paper and the ADS algorithm in literature Suo et al. (2024c) is analyzed. Both algorithms aim to select the set of malicious information that have the greatest impact on the objective function (6). Therefore, ideally, their performance is consistent, meaning that $\mathcal{A}_{i,\,k}=\cup^{m_{i}}_{g_{i}=1}{\mathcal{A}_{i,\,k,\,g_{i}}}$.
@@
For $g_{i}\in\{1,\,...,\,m_{i}\}$, $l\in\{0,\,1,\,...,\,q_{i}\}$, define $\delta_{l,\,g_{i}}$ as $\delta_{l,\,g_{i}}=\sum_{k=1}^{T}{(}\frac{1}{m_{i}}f_{k}(\mathcal{A}_{i,\,k}^{*})-f_{k}(\mathcal{A}_{i,\,k,\,g_{i}}^{(l_{g_{i}})}))$ , where $\mathcal{A}_{i,\,k,\,g_{i}}^{(l_{g_{i}})}$ represents the set of attacked sensors after the $l_{g_{i}}$-th selection of the $g_{i}$-th subset at moment $k$. Then, the relationship in literature Suo et al. (2024c) can be transformed as $\sum^{m_{i}}_{g_{i}=1}\delta_{l+1,\,g_{i}}-(1-\frac{1}{q_{i}})^{l+1}\sum^{m_{i}}_{g_{i}=1}\delta_{0,\,g_{i}}\leq\frac{m_{i}}{q_{i}}\sum_{j=1}^{l+1}(1-\frac{1}{q_{i}})^{l+1-j}B_{i}^{(j)}$. The parameter $B_{i}^{(l)}$ is described in detail in the proof. At this time, the Lemma 3.2 in literatureSuo et al. (2024c) is extended to the distributed case.
@@
$\frac{1}{m_{i}}\sum^{m_{i}}_{g_{i}=1}\delta_{l,g_{i}}$ $\leq$ $\frac{1}{m_{i}}\sum^{T}_{k=1}(\sum^{m_{i}}_{g_{i}=1}(f_{k}(\mathcal{A}_{i,\,k,\,g_{i}}^{*})-f_{k}(\mathcal{A}_{i,\,k,\,g_{i}}^{(l_{g_{i}})})))$ (28)
@@
$\frac{1-q_{i}}{m_{i}}\sum^{m_{i}}_{g_{i}=1}\delta_{l,\,g_{i}}\leq-\frac{q_{i}}{m_{i}}\left(\sum^{m_{i}}_{g_{i}=1}\delta_{l+1,\,g_{i}}\right)+B^{(l+1)}_{i}$, (29)
@@
$B^{(l)}_{i}=\sum_{j=1}^{q_{i}}\sum_{k=1}^{T}\left(\frac{1}{m_{i}}\sum^{m_{i}}_{g_{i}=1}G_{k,\,g_{i}}^{(l_{g_{i}})}p_{g_{i},\,k}^{(l_{g_{i}})}-G_{kj_{kl}^{*}}^{(l)}\right),$ (30)
@@
$\sum^{m_{i}}_{g_{i}=1}\delta_{l+1,\,g_{i}}-\left(1-\frac{1}{q_{i}}\right)\sum^{m_{i}}_{g_{i}=1}\delta_{l,\,g_{i}}\leq\frac{m_{i}}{q_{i}}B_{i}^{(l+1)}$ (31)
@@
For the dynamic attack strategy in the Definition 1, the proposed D-ADS algorithm can ensure that the theoretical lower bound of the average optimization rate of malicious information selection for any subset is 11/e1-1/e, and the error expectation is bounded.
@@
To prove that the theoretical lower bound of the average optimization rate over the entire time period is 11/e1-1/e, we need to show that the expectation of the error
@@
is bounded, where $f_{k}\left(\mathcal{A}_{i,\,k,\,g_{i}}\right)$ represents the submodular function value for the selected malicious information set in subset $g_{i}$ at time $k$.
Essentially, the expectation of the error represents the expected difference between the objective function value of the selected malicious information set and 11/e1-1/e times the optimal malicious information set’s objective function value. Therefore, the larger the objective function value of a subset, the greater the upper bound of the error expectation. Consequently, the subset with the largest objective function value has the highest upper bound for error expectation, as shown in equation (B)
@@
Based on the Lemma 3 from literature Matsuoka et al. (2021) and the definition $\delta_{l}=\sum_{k=1}^{T}\left(f_{k}(\mathcal{A}_{i,\,k}^{*})-f_{k}(\mathcal{A}_{i,\,k}^{(l)})\right)$ for $l=1:q_{i}$, we obtain that, $\sum_{k=1}^{T}{f_{k}\left(\mathcal{A}_{i,\,k}^{*}\right)}-\sum_{k=1}^{T}{f_{k}\left(\mathcal{A}_{i,\,k}\right)}+(1-(1-\frac{1}{q_{i}})^{q_{i}})\sum_{k=1}^{T}{(f_{k}(\mathcal{A}_{i,\,k}^{(0)})-f_{k}(\mathcal{A}_{i,\,k}^{*}))}\leq\delta_{q_{i}}-(1-\frac{1}{q_{i}})^{q_{i}}\delta_{0}\leq\sum^{m_{i}}_{g_{i}=1}\delta_{l,\,g_{i}}-(1-\frac{1}{q_{i}})^{q_{i}}\sum^{m_{i}}_{g_{i}=1}\delta_{0,\,g_{i}}$, where the inequality $\delta_{q_{i}}\leq\sum^{m_{i}}_{g_{i}=1}\delta_{l,\,g_{i}}$ follows from the diminishing marginal returns property of the submodular function, and the initial condition satisfies $\delta_{0}=\sum_{g_{i}=1}^{m_{i}}\delta_{0,\,g_{i}}$. Combining equation (32) with the above relation, equation (B) can be rewritten as
$(1-\frac{1}{e})\sum_{k=1}^{T}{\frac{1}{m_{i}}f_{k}\left(\mathcal{A}_{i,\,k}^{*}\right)}-\sum_{k=1}^{T}{\max_{g_{i}}f_{k}(\mathcal{A}_{i,\,k,\,g_{i}})}\leq\frac{m_{i}}{q_{i}}\sum_{j=1}^{q_{i}}(1-\frac{1}{q_{i}})^{q_{i}-j}\frac{B_{i}^{(j)}}{m_{i}}.$ (35)
Thus, to prove that equation (B) is bounded, it suffices to show that $\mathbb{E}[B_{i}^{(l)}/m_{i}]$ is bounded
$\mathbb{E}\left[\frac{B_{i}^{(l)}}{m_{i}}\right]$ $=$ $\frac{1}{m_{i}}\sum_{j=1}^{q_{i}}\mathbb{E}\left[\sum_{k=1}^{T}\left(\frac{1}{m_{i}}\sum^{m_{i}}_{g_{i}=1}G_{k,\,g_{i}}^{(l_{g_{i}})}p_{g_{i},\,k}^{(l_{g_{i}})}-G_{kj_{kl}^{*}}^{\left(l\right)}\right)\right]$
@@
$\mathbb{E}\left[\left(1-\frac{1}{e}\right)\sum_{k=1}^{T}{\frac{1}{m_{i}}f_{k}\left(\mathcal{A}_{i,\,k}^{*}\right)}-\sum_{k=1}^{T}{f_{k}(\mathcal{A}_{i,\,k,\,g_{i}})}\right]\leq\tilde{\mathcal{O}}\left(\frac{q_{i}}{m_{i}}\sqrt{3T+2\varDelta_{T}}\right),$ (37)
@@
Acknowledgments
The authors note that this paper has also been made available as a preprint on arXiv (Suo et al., 2025).